# Randomization Tests for Conditional Group Symmetry

### v1

@@ 7.3 Large Hadron Collider dijet events
Table 1: Rejection rates (over $1000$ repetitions) for tests for $\mathrm{SO}(2)$-equivariance on LHC dijet data samples of size $n=100$.
@@
Comparison set BS-FS BS-SK RN-FS RN-SK
Original data Reused $0.000$ $0.002$ $0.000$ $0.000$
Independent $0.002$ $0.003$ $0.005$ $0.004$
Shuffled data Reused $0.818$ $0.836$ $0.922$ $0.905$
Independent $0.838$ $0.815$ $0.901$ $0.889$
Our first application demonstrates how our tests could be used to verify symmetry presumed by theory and to detect possible instances of symmetry breaking for scientific discovery. The Large Hadron Collider (LHC) Olympics 2020 dataset (Kasieczka et al., 2021; Kasieczka, Nachman and Shih, 2019) consists of 1.1 million simulated dijet events generated by PYTHIA (Bierlich et al., 2022), a widely-used Monte Carlo simulator for high-energy physics processes. A dijet is two jets of constituent particles produced from the decay of high-energy quarks following a particle collision. The transverse momentum, polar angle $\phi$ and pseudorapidity $\eta$ for up to 200 constituents were recorded for each jet; the Cartesian momentum in the transverse plane for each constituent can be computed from these values as the pair $(p_{1}=p_{\text{T}}\cos(\phi),p_{2}=p_{\text{T}}\sin(\phi))$. The leading constituents of a jet are the particles with the largest transverse momenta in any direction.
By conservation of angular momentum, the Cartesian momenta of one leading constituent, $Y=(p_{1,Y},p_{2,Y})$, should be equivariant with respect to $\mathrm{SO}(2)$ rotations of the other, $X=(p_{1,X},p_{2,X})$. We conduct tests for $\mathrm{SO}(2)$-equivariance on random samples of size $n=100$ drawn from the full data. The estimated rejection rates are shown in Table 1. Regardless of whether a randomized comparison set is reused or not, we find that all tests achieve a rejection rate of no more than $0.005$, suggesting the presence of $\mathrm{SO}(2)$-equivariance, which is consistent with the theory. As a check for power, we also consider tests for $\mathrm{SO}(2)$-equivariance on samples of the same size where the second leading constituents were permuted between observations, breaking the dependence and thus the conditional equivariance. The tests all reject with a power greater than $0.8$, which suggests the absence of symmetry in this setting.
Figure 18 in Section D.7 visualizes the transverse momentum of the leading constituents. It can be observed that the momenta $X$ of the first leading constituents appear to be distributed on a ring centered at the origin, suggesting that the distribution of $X$ may be marginally invariant with respect to the action of $\mathrm{SO}(2)$. As per Property E7, equivariance could then be tested via a test for joint invariance under the action of the group of paired rotations $\mathbf{G}_{0}=\{(g_{1},g_{2})\in\mathrm{SO}(2)\times\mathrm{SO}(2):g_{1}=g_{2}\}$. We conduct a test for invariance with respect to this subgroup on samples of size $n=25$. Furthermore, as the presence of unexpected symmetry would also be of scientific interest, we also test for invariance under the action of the full $\mathbf{G}_{1}=\mathrm{SO}(2)\times\mathrm{SO}(2)$ group and of $\mathbf{G}_{2}=\mathrm{SO}(4)$. The results are shown in Table 2. We see that our tests detect $\mathbf{G}_{0}$-invariance and reject $\mathbf{G}_{1}$- and $\mathbf{G}_{2}$-invariance, which is again consistent with the Standard Model.
Test
Group to test Comparison set BS-FS BS-SK RN-FS RN-SK
[…]
Independent $0.044$ $0.048$ $0.059$ $0.040$
[…]
Independent $0.425$ $0.838$ $0.946$ $0.970$
$\mathrm{SO}(4)$ Reused $0.615$ $0.915$ $0.983$ $0.976$
Independent $0.611$ $0.917$ $0.982$ $0.977$
@@ 7.4 Lorentz symmetry of quark decays
We consider a second particle physics application based on the Top Quark Tagging Reference dataset (Kasieczka et al., 2019), which also consists of jet events simulated by PYTHIA. The original dataset was constructed for the task of classifying jet events as having decayed from a top quark or not and consists of a training, validation, and test set. The four-momenta $x=(E,p_{1},p_{2},p_{3})$ of up to 200 jet constituents ordered by transverse momentum (i.e., highest magnitude momentum in the plane perpendicular to the particle beam line) are recorded for each event. We only use the test set, which contains 404,000 simulated events, and ignore the binary label that indicates whether the jet decayed from a top quark or not. According to the Standard Model, the four-momenta of the constituents in a jet should be equivariant with respect to the Lorentz group $\mathrm{SO}^{+}(1,3)$, which was described in Examples 4 and 6, and preserves the quadratic form $Q(x)=E^{2}-p_{1}^{2}-p_{2}^{2}-p_{3}^{2}$.
We take $X$ to be the four-momentum of the leading constituent in each jet, and $Y$ to be that of the second leading constituent (as in Yang et al., 2023). We conduct tests for $\mathrm{SO}^{+}(1,3)$-equivariance of $P_{Y\mid X}$ based on samples of size $n=100$ using the maximal invariant $Q(x)$. The estimated rejection rates are shown in Table 3. The tests all reject $\mathrm{SO}^{+}(1,3)$-equivariance at a rate of less than $0.01$, indicating symmetry consistent with the Standard Model. Like in the LHC dijet experiment, we also conduct tests on samples where the second leading constituents are permuted across the data. In this setting, the conditional randomization tests are able to detect the breaking of $\mathrm{SO}^{+}(1,3)$-equivariance with reasonable power. The baseline permutation tests are not as powerful in this setting.
@@
Comparison set BS-FS BS-SK RN-FS RN-SK
Original data Reused $0.000$ $0.001$ $0.002$ $0.004$
Independent $0.000$ $0.001$ $0.006$ $0.009$
Shuffled data Reused $0.281$ $0.195$ $0.780$ $0.752$
Independent $0.281$ $0.195$ $0.686$ $0.650$
7.5 Training an invariant MNIST classifier
As our final application, we illustrate how our tests for symmetry of a conditional distribution $P_{Y|X}$ can be adapted as tests for symmetry of a function $f:\mathbf{X}\to\mathbf{Y}$ by taking $P_{Y|X}=\delta_{f(X)}$. In light of this, our tests could be used as, for example, validation tools for assessing whether a model has learned symmetry after being trained on data. These test-based tools would only require data in the form of input and predicted output pairs. To demonstrate this utility, we consider a setting where a MNIST classifier is trained to be invariant with respect to rotations of MNIST images via data augmentation (Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022). Data augmentation aims to desensitize the model to certain transformations by randomly transforming the training data. Therefore, unlike methods that hardcode symmetry as inductive biases, it may not always be obvious whether a model has learned symmetry until after it is applied in practice.
The MNIST dataset is a collection of $28\times 28$ pixel images of handwritten digits. The training set consists of $20000$ images; the test set consists of $45000$ augmented images, from which we uniformly sample to use in our tests. We train two identical LeNet convolutional neural networks for predicting the digit in the image, where one model is trained on images augmented with rotations and the other without. Before and during training, we pause the training process after every three epochs to test the invariance properties of each model. To generate the data used in each test, a sample of $n=100$ augmented images is first sampled from the test set and then fed into the model under examination. The data $Y$ are then taken as the digit probabilities predicted by the model, and so each observation is a vector on a simplex of dimension $d=10$ (or $d=9$ in the setting where the $9$ images is removed from the data). In our tests, we take the maximal invariant to be the unaugmented image to avoid complications with orienting images.
Figure 8: CRT rejection rates for image rotation-invariance of two MNIST models during training, where one is trained with data augmentation and the other without. The “R” prefix in the test label denotes reuse of a comparison set; the “I” prefix denotes independent comparison sets. The top row shows results for when the models are trained on the full MNIST data, and the bottom row shows results for when images with true label $9$ are removed from the data.
Here, we used the estimated test rejection rate as a measure of model invariance. It may be tempting to determine whether symmetry has been learned at a particular time point by whether a single test rejects or not, but conducting a suite of tests at multiple points during training would be subject to multiple testing considerations. A potential direction for future work is anytime-valid tests for conditional symmetry, which would allow for the continual evaluation of symmetry-learning based on a single test at each time point during training. Anytime-valid tests for marginal group invariance have recently appeared in the literature (Pérez-Ortiz et al., 2024; Lardy and Pérez-Ortiz, 2024).
Appendix A Background on groups
@@
A group $\mathbf{G}$ is a set with a binary operation $\cdot$ that satisfies the associativity, identity, and inverse axioms. We denote the identity element by id. For notational convenience, we write $g_{1}g_{2}=g_{1}\cdot g_{2}$ for $g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ is said to be measurable if the group operations $g\mapsto g^{-1}$ and $(g_{1},g_{2})\mapsto g_{1}g_{2}$ are $\mathbf{S}_{\mathbf{G}}$-measurable, where $\mathbf{S}_{\mathbf{G}}$ is a $\sigma$-algebra of subsets of $\mathbf{G}$. A group action is a function $\Phi\colon\mathbf{G}\times\mathbf{X}\to\mathbf{X}$ that satisfies $\Phi(\textrm{id},x)=x$ and $\Phi(g_{1}g_{2},x)=\Phi(g_{1},\Phi(g_{2},x))$ for all $x\in\mathbf{X},g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is measurable relative to $\mathbf{S}_{\mathbf{G}}\otimes\mathbf{S}_{\mathbf{X}}$ and $\mathbf{S}_{\mathbf{X}}$. Throughout, all actions are assumed to be continuous (and therefore measurable), and for convenience we simply say that $\mathbf{G}$ acts on $\mathbf{X}$, writing $gx=\Phi(g,x)$. This standard notation hides the particular action $\Phi$ associated with $gx$, and it is important not to lose track of that fact.
@@ A.1 Proper group actions
In the analysis of probabilistic aspects of group actions, measurability issues can arise without regularity conditions. The key regularity condition that we assume is that the group action is proper: there exists a strictly positive measurable function $h\colon\mathbf{X}\to\mathbb{R}_{+}$ such that $\int_{\mathbf{G}}h(gx)\lambda(dg)<\infty$ for $x\in\mathbf{X}$ (Kallenberg, 2007). This definition of proper group action is a weaker, non-topological version of the definition commonly used in previous work in the statistics literature (e.g., Eaton, 1989; Wijsman, 1990; McCormack and Hoff, 2023), and only requires the existence of Haar measure. The common topological version defines $(g,x)\mapsto(gx,x)$ as a proper map, i.e., the inverse image of each compact set in $\mathbf{X}\times\mathbf{X}$ is a compact set in $\mathbf{G}\times\mathbf{X}$. That definition implies the one we use; see Kallenberg (2007) for details.
@@
To state the proposition, let $\nu$ be any bounded measure on $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ and let $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ be the completion of $\mathbf{S}_{\mathbf{X}}$ to include all subsets of $\nu$-null sets, and denote by $\bar{\nu}$ the extension of $\nu$ to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ (see, e.g. Çinlar, 2011, Proposition 1.3.10). All statements of $\bar{\nu}$-measurability in the following proposition are with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, so that a set $A\subseteq\mathbf{X}$ is $\bar{\nu}$-measurable if $A\in\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$. Moreover, a function defined by a particular property is $\bar{\nu}$-measurable if it is measurable in the usual sense with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, and if the defining property holds with the possible exception of a $\bar{\nu}$-null set. Such a function would also be $\bar{\rho}$-measurable for any measure $\rho\ll\nu$.
@@
    There exists a $\bar{\nu}$-measurable orbit selector $\rho\colon\mathbf{X}\to[\mathbf{X}]$, which is a maximal invariant statistic, and it induces a $\bar{\nu}$-measurable cross-section $[\mathbf{X}]=\rho(\mathbf{X})$.
@@
    For a fixed $\bar{\nu}$-measurable orbit selector $\rho$, there exists a unique $\bar{\nu}$-measurable inversion probability kernel $P_{\tilde{\tau}|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ with the following properties:
@@
      $P_{\tilde{\tau}|X}$ is $\mathbf{G}$-equivariant: For all $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{G}}$, $P_{\tilde{\tau}|X}(B|gx)=P_{\tilde{\tau}|X}(g^{-1}B|x)$.
@@
      For each $x\in\mathbf{X}$, $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(x))$ is normalized Haar measure on the stabilizer subgroup $\mathbf{G}_{\rho(x)}$.
@@
      For each $x\in\mathbf{X}$, if $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$, then $\tilde{\tau}\rho(x)=x$ with probability one.
@@
      If there is a $\bar{\nu}$-measurable representative inversion $\tau\colon\mathbf{X}\to\mathbf{G}$ associated with $\rho$ such that it satisfies $\tau(x)\rho(x)=x$ and $\tau(gx)=g\tau(x)$ for each $x\in\mathbf{X},g\in\mathbf{G}$, then $P_{\tilde{\tau}|X}^{\prime}(B|x)=P_{\tilde{\tau}|X}(\tau(x)^{-1}B|\rho(x))$ is an equivalent inversion kernel. In particular, this holds when the action of $\mathbf{G}$ on $\mathbf{X}$ is free, in which case $\mathbf{G}_{\rho(x)}=\{\textrm{id}\}$ and the inversion kernel is $\delta_{\tau(x)}$.
The measurability of the canonical projection is a result from functional analysis; see Eaton (1989, Theorem 5.4) for an extended statement and references. One implication is that $\pi$ generates the invariant $\sigma$-algebra on $\mathbf{X}$, so that every invariant function can be written as a measurable function of it. Items 2–3c follow directly from results of Kallenberg (2011, 2017) on the existence of universally measurable versions of $\rho$ and $P_{\tilde{\tau}|X}$. Item 3d follows from 3a and 3b.
@@
Let $\mathbf{G}$ act properly on $\mathbf{X}$, $P_{\Gamma|[X]}$ be as defined in (9), and $P_{X}$ be any probability measure on $\mathbf{X}$. If $X\sim P_{X}$ then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$.
@@
For any measurable $f\colon[\mathbf{X}]\times\mathbf{X}\to\mathbb{R}$, by 3(c) of Proposition A.1, standard disintegration of measure, and Fubini’s theorem,
$\int$ $f(\rho(x),x)P_{X}(dx)=\int f(\rho(x),\tilde{\tau}\rho(x))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f([x],\tilde{\tau}[x])P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x^{\prime})P_{X\mid[X]}(dx^{\prime}\mid[x])(\rho_{*}P_{X})(d[x])$
$=\int f([x],\gamma[x])P_{\Gamma|[X]}(d\gamma\mid[x])(\rho_{*}P_{X})(d[x])$
$=\int f(\rho(x),\gamma\rho(x))P_{\Gamma|[X]}(d\gamma\mid\rho(x))P_{X}(dx)\;.$
Appendix B Proofs
B.1 Proof of Theorem 2.2
The proof of Theorem 2.2 relies on two auxiliary results. The first is a characterization of distributional invariance, which is well-studied in probability (Kallenberg, 2005). See Eaton (1989) for an accessible treatment of the basic results and some of their applications in classical statistics. We summarize two characterizations of invariance that are needed below.

### v2

@@ 6.3 Large Hadron Collider dijet events
Our first application demonstrates how our tests could be used to verify symmetry presumed by theory and to detect possible instances of symmetry breaking for scientific discovery. The Large Hadron Collider (LHC) Olympics 2020 dataset (Kasieczka et al., 2021; Kasieczka, Nachman and Shih, 2019) consists of 1.1 million simulated dijet events generated by PYTHIA (Bierlich et al., 2022), a widely-used Monte Carlo simulator for high-energy physics processes. A dijet is two jets of constituent particles produced from the decay of high-energy quarks following a particle collision. The transverse momentum $p_{T}$ and polar angle $\phi$ for up to 200 constituents were recorded for each jet; the Cartesian momentum in the transverse plane for each constituent can be computed from these values as the pair $(p_{1}=p_{\text{T}}\cos(\phi),p_{2}=p_{\text{T}}\sin(\phi))$. The leading constituents of a jet are the particles with the largest transverse momenta in any direction.
Table 1: Rejection rates (over $1000$ repetitions) for tests for $\mathrm{SO}(2)$-equivariance on LHC dijet data samples of size $n=25$.
@@
BS-FS BS-SK RN-FS RN-SK
Original data $0.003$ $0.001$ $0.000$ $0.000$
Modified data $0.694$ $0.665$ $0.796$ $0.817$
By conservation of angular momentum, the Cartesian momenta of one leading constituent, $Y=(p_{1,Y},p_{2,Y})$, should be equivariant with respect to $\mathrm{SO}(2)$ rotations of the other, $X=(p_{1,X},p_{2,X})$. We conduct tests for $\mathrm{SO}(2)$-equivariance on random samples of size $n=25$ drawn from the full data. The estimated rejection rates based on $B=1000$ randomizations are shown in Table 1. We find that all tests achieve an average rejection rate of no more than $0.005$, suggesting the presence of $\mathrm{SO}(2)$-equivariance, which is consistent with physical theory. As a check for power, we also consider tests for $\mathrm{SO}(2)$-equivariance on samples of the same size, where the momenta of $Y$ are replaced by $Y^{\prime}=(|p_{1,Y}|,p_{2,Y})$, breaking the conditional equivariance. The tests all reject with a power greater than $0.65$, which suggests the absence of symmetry in this setting. Here, note that an equivalent test for $\mathrm{SO}(2)$-equivariance is a test for joint $\mathrm{SO}(2)$-invariance. In Section E.8, we demonstrate what a test for joint invariance on these data would look like.
@@ 6.4 Lorentz symmetry of quark decays
We consider a second particle physics application based on the Top Quark Tagging Reference dataset (Kasieczka et al., 2019), which also consists of jet events simulated by PYTHIA. The original dataset was constructed for the task of classifying jet events as having decayed from a top quark or not and consists of a training, validation, and test set. The four-momenta $x=(E,p_{1},p_{2},p_{3})$ of up to 200 jet constituents ordered by transverse momentum (i.e., highest magnitude momentum in the plane perpendicular to the particle beam line) are recorded for each event. We only use the test set, which contains 404,000 simulated events, and ignore the binary label that indicates whether the jet decayed from a top quark or not. According to the Standard Model, the four-momenta of the constituents in a jet should be equivariant with respect to the Lorentz group $\mathrm{SO}^{+}(1,3)$, which was described in Examples 3 and 5, and preserves the quadratic form $Q(x)=E^{2}-p_{1}^{2}-p_{2}^{2}-p_{3}^{2}$.
We take $X$ to be the four-momentum of the leading constituent in each jet, and $Y$ to be that of the second leading constituent (as in Yang et al., 2023). We conduct tests for $\mathrm{SO}^{+}(1,3)$-equivariance of $P_{Y\mid X}$ based on samples of size $n=100$ using the maximal invariant $Q(x)$. The estimated rejection rates based on $B=1000$ randomizations are shown in Table 2. The tests all reject $\mathrm{SO}^{+}(1,3)$-equivariance at a rate of less than $0.01$, indicating symmetry consistent with the Standard Model. We check for power by permuting the second leading constituents, breaking the dependence and thus the conditional equivariance. In this setting, the conditional randomization tests are able to detect the breaking of $\mathrm{SO}^{+}(1,3)$-equivariance with reasonable power. The baseline permutation tests are not as powerful in this setting, demonstrating the statistical benefit of using the group structure in the randomization scheme.
@@
BS-FS BS-SK RN-FS RN-SK
Original data $0.000$ $0.002$ $0.001$ $0.002$
Shuffled data $0.292$ $0.175$ $0.812$ $0.768$
Appendix A Additional background on groups
A group $\mathbf{G}$ is a set with a binary operation $\cdot$ that satisfies the associativity, identity, and inverse axioms. We denote the identity element by id. For notational convenience, we write $g_{1}g_{2}=g_{1}\cdot g_{2}$ for $g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ is said to be measurable if the group operations $g\mapsto g^{-1}$ and $(g_{1},g_{2})\mapsto g_{1}g_{2}$ are $\mathbf{S}_{\mathbf{G}}$-measurable, where $\mathbf{S}_{\mathbf{G}}$ is a $\sigma$-algebra of subsets of $\mathbf{G}$. For $A\subseteq\mathbf{G}$ and $g\in\mathbf{G}$, we write $gA=\{gh:h\in A\}$ and $Ag=\{hg:h\in A\}$. A measure $\nu$ on $\mathbf{G}$ is said to be left-invariant if $\nu(gA)=\nu(A)$ for all $A\in\mathbf{S}_{\mathbf{G}}$, and right-invariant if $\nu(Ag)=\nu(A)$. When $\mathbf{G}$ is lcscH, there exist left- and right-invariant $\sigma$-finite measures $\lambda_{\mathbf{G}}$ and $\tilde{\lambda}_{\mathbf{G}}$, known as left- and right-Haar measures, respectively, that are unique up to scaling (Folland, 2016, Ch. 2.2).
@@ Appendix A Additional background on groups
A group action is a function $\Phi\colon\mathbf{G}\times\mathbf{X}\to\mathbf{X}$ that satisfies $\Phi(\textrm{id},x)=x$ and $\Phi(g_{1}g_{2},x)=\Phi(g_{1},\Phi(g_{2},x))$ for all $x\in\mathbf{X},g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is measurable relative to $\mathbf{S}_{\mathbf{G}}\otimes\mathbf{S}_{\mathbf{X}}$ and $\mathbf{S}_{\mathbf{X}}$. Throughout, all actions are assumed to be continuous (and therefore measurable), and for convenience we simply say that $\mathbf{G}$ acts on $\mathbf{X}$, writing $gx=\Phi(g,x)$. This standard notation hides the particular action $\Phi$ associated with $gx$, and it is important not to lose track of that fact.
@@ A.1 Proper group actions
In the analysis of probabilistic aspects of group actions, measurability issues can arise without regularity conditions. The key regularity condition that we assume is that the group action is proper: there exists a strictly positive measurable function $h\colon\mathbf{X}\to\mathbb{R}_{+}$ such that $\int_{\mathbf{G}}h(gx)\lambda(dg)<\infty$ for every $x\in\mathbf{X}$ (Kallenberg, 2007). This definition of proper group action is a non-topological version of the definition commonly used in previous work in the statistics literature (e.g., Eaton, 1989; Wijsman, 1990; McCormack and Hoff, 2023), and only requires the existence of Haar measure. The common topological version defines $(g,x)\mapsto(gx,x)$ as a proper map, i.e., the inverse image of each compact set in $\mathbf{X}\times\mathbf{X}$ is a compact set in $\mathbf{G}\times\mathbf{X}$. That definition implies the one we use; see Kallenberg (2007) for details.
@@
To state the proposition, let $\nu$ be any bounded measure on $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ and let $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ be the completion of $\mathbf{S}_{\mathbf{X}}$ to include all subsets of $\nu$-null sets, and denote by $\bar{\nu}$ the extension of $\nu$ to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ (see, e.g. Çinlar, 2011, Proposition 1.3.10). All statements of $\bar{\nu}$-measurability in the following proposition are with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, so that a set $A\subseteq\mathbf{X}$ is $\bar{\nu}$-measurable if $A\in\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$. Moreover, a function defined by a particular property is $\bar{\nu}$-measurable if it is measurable in the usual sense with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, and if the defining property holds with the possible exception of a $\bar{\nu}$-null set. Such a function would also be $\bar{\upsilon}$-measurable for any measure $\upsilon\ll\nu$.
@@
    There exists a $\bar{\nu}$-measurable orbit selector $\rho\colon\mathbf{X}\to\mathbf{R}$, which is a maximal invariant statistic, and it induces a $\bar{\nu}$-measurable cross-section $\mathbf{R}$.
@@
    For a fixed $\bar{\nu}$-measurable orbit selector $\rho$, there exists a unique $\bar{\nu}$-measurable inversion probability kernel $P_{\Psi|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ with the following properties:
@@
      $P_{\Psi|X}$ is $\mathbf{G}$-equivariant: For all $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{G}}$, $P_{\Psi|X}(B|gx)=P_{\Psi|X}(g^{-1}B|x)$.
@@
      For each $x\in\mathbf{X}$, $P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(x))$ is normalized Haar measure on the stabilizer subgroup $\mathbf{G}_{\rho(x)}$.
@@
      For each $x\in\mathbf{X}$, if $\Psi|X=x\sim P_{\Psi|X}$, then $\Psi\rho(x)=x$ with probability one.
@@
      If there is a $\bar{\nu}$-measurable representative inversion ${\psi^{\circ}}\colon\mathbf{X}\to\mathbf{G}$ associated with $\rho$ such that it satisfies ${\psi^{\circ}}(x)\rho(x)=x$ and ${\psi^{\circ}}(gx)=g{\psi^{\circ}}(x)$ for each $x\in\mathbf{X},g\in\mathbf{G}$, then $P_{\Psi|X}^{\prime}(B|x)=P_{\Psi|X}({\psi^{\circ}}(x)^{-1}B|\rho(x))$ is an equivalent inversion kernel. In particular, this holds when the action of $\mathbf{G}$ on $\mathbf{X}$ is free, in which case $\mathbf{G}_{\rho(x)}=\{\textrm{id}\}$ and the inversion kernel is $\delta_{{\psi^{\circ}}(x)}$.
The measurability of the canonical projection is a result from functional analysis; see Eaton (1989, Theorem 5.4) for an extended statement and references. One implication is that $\pi$ generates the invariant $\sigma$-algebra on $\mathbf{X}$, so that every invariant function can be written as a measurable function of it. Items 2–3c follow directly from results of Kallenberg (2011, 2017) on the existence of universally measurable versions of $\rho$ and $P_{\Psi|X}$. Item 3d follows from 3a and 3b.
@@
Let $\mathbf{G}$ act properly on $\mathbf{X}$, $P_{\Gamma|R}$ be as defined in (8), and $P_{X}$ be any probability measure on $\mathbf{X}$. If $X\sim P_{X}$, then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$.
@@
For any measurable $f\colon\mathbf{R}\times\mathbf{X}\to\mathbb{R}$, by 3(c) of Proposition A.1, standard disintegration of measure, and Fubini’s theorem,
$\int f(\rho(x),x)P_{X}(dx)$ $=\int f(\rho(x),\Psi\rho(x))P_{\Psi|X}(d\Psi\mid x)P_{X}(dx)$
$=\int f(r,\Psi r)P_{\Psi|X}(d\Psi\mid x^{\prime})P_{X\mid R}(dx^{\prime}\mid r)(\rho_{*}P_{X})(dr)$
$=\int f(r,\gamma r)P_{\Gamma|R}(d\gamma\mid r)(\rho_{*}P_{X})(dr)$
$=\int f(\rho(x),\gamma\rho(x))P_{\Gamma|R}(d\gamma\mid\rho(x))P_{X}(dx)\;.$
Appendix B Further background on testing for distributional invariance
As noted in Section 2.2, the literature appears to be missing a proof of the conditional level for the randomization test in the most general setting of a compact group acting properly, with possible non-trivial stabilizer subgroups.
For simplicity, we assume that the probability of ties is zero. If that is not the case (i.e., if $T$ is supported on a discrete set), then a modified version with randomized tie-breaking yields a level-$\alpha$ test (e.g., Hemerik and Goeman, 2018).
Theorem B.1.
Let $\mathbf{G}$ be a compact group that acts on $\mathbf{X}$. Let $M$ be any maximal invariant. Assume that $X_{1:n}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P$. For test statistic $T$ and the randomization test in Algorithm 1, assume that $\mathbb{P}\{T(X^{(b)}_{1:n})=T(X^{(b^{\prime})}_{1:n})\}=0$ for $b\neq b^{\prime}$. For any fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, if $P$ is $\mathbf{G}$-invariant then
$\mathbb{P}\left(p_{B}\leq\alpha\mid M(X)_{1:n}\right)=\frac{\lfloor\alpha(B+1)\rfloor}{B+1}\leq\alpha\;,\quad P\text{-a.s.}$
The high-level idea of the proof is as follows. It is known that any maximal invariant is a sufficient statistic for $\mathcal{P}^{\circ}(\mathbf{X})$, the set of $\mathbf{G}$-invariant probability distributions on $\mathbf{X}$ (Farrell, 1962; Dawid, 1985; Bloem-Reddy and Teh, 2020). The use of sufficiency greatly simplifies the subsequent technical details, allowing for the application of standard results for Monte Carlo tests, and also exposing randomization tests for invariance to be co-sufficient sampling methods (Barber and Janson, 2022).
The proof of Theorem B.1 relies on two auxiliary results. The first is a characterization of distributional invariance, which is well-studied in probability (Kallenberg, 2005). See Eaton (1989) for an accessible treatment of the basic results and some of their applications in classical statistics. We summarize two characterizations of invariance that are needed below.